# EFX Exists for Three Types of Agents

### v1

@@ 48                         $U\leftarrow U\cup(X_{c_{i}}\setminus P)$
Algorithm 1 Multi-Round-Competition
@@
    Finally, if $P=X_{a_{i}}$ for some $a_{i}\in\mathbb{A}\setminus a_{1}$, then give $P$ to agent $b_{1}$ and $\{T_{b_{1}}^{a_{1}}\cup g\}$ to agent $a_{i}$. Resolve internal envies in $\mathbb{B}$ or $\mathbb{A}$, if any, using Lemma 4. We now have a EFX allocation $Y$ such that agent $a_{1}$ envies $b_{1}$. Therefore, by applying Lemma 8 on $Y$, we get an EFX allocation $Z$ such that $\phi(Z)>\phi(X)$.
@@
A virtual bundle appears in $O_{b}$ when Case 3Case 3a holds i.e. $\max_{b}O_{b}=X_{c_{i}}$ for some $i$ and $\max_{c}O_{c}=X_{b_{j}}$ for some $j\neq 1$, or Case 3(Case 3b)Case 3b(ii) holds, when $P=Q=X_{c_{i}}$ and either $b_{1}$ has a smaller winning subset than $c_{i}$ i.e. $|P_{b}|<|P_{c}|$ or both have $X_{c_{i}}$ as the winning subset i.e. $P_{b}=P_{c}=X_{c_{i}}$. When $P=\max_{b}O_{b}$ is a virtual bundle, then for any non-virtual i.e. real bundle $S\in O_{b}$, $P>_{b}S$ and hence $b_{1}$ does not envy the agent holding $S$. For any virtual bundle $S=X_{c_{i}}\setminus h\in O_{b}$, $S\neq P$, $P>_{b}X_{c_{i}}\setminus h$ for minimum valued item $h\in X_{c_{i}}$ according to $v_{b}$. Hence $b_{1}$ might envy the real bundle $X_{c_{i}}$ but it is not a strong envy. Since $b_{j}$ and $b_{1}$ have the same valuation $v_{b}$, the lemma holds for any $b_{j}\in B$.∎
To prove the correctness of Case 3, one needs to show that Invariant 1 is maintained as long as Algorithm 1 remains in Case 3, and that the algorithm exits from Case 3 in a finite number of iterations.
@@ 6 Agent $a_{1}$ is Not a Source in $E_{X}$
since $a_{1}$ is not a source, either $b_{1}$ or $c_{1}$ must envy $a_{1}$ by Proposition 4. Without loss of generality, assume $b_{1}$ envies $a_{1}$.
  1. 1
    Agent $a_{1}$ envies some agent $b_{j}\in\mathbb{B}$: Then, $\langle(a_{1},b_{j}),(b_{1},a_{1})\rangle$ is a $\phi$-improving exchange. Therefore, using Lemma 3 we get a new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
  2. 2
    Agent $a_{1}$ envies some $c_{j}\in\mathbb{C}$: If $c_{1}$ envies $a_{1}$, then $\langle(a_{1},c_{j}),(c_{1},a_{1})\rangle$ is a $\phi$-improving exchange. Thus we are done. Consider the case when $c_{1}$ does not envy $a_{1}$. We know that some leading agent must $g$-champion $b_{1}$. We consider all possible cases based on who $g$-champions $b_{1}$.
    1. 2.1
@@
    2. 2.2
      Agent $a_{1}$ $g$-champions $b_{1}$: Then $\langle(a_{1},b_{1}),(b_{1},a_{1})\rangle$ is a $\phi$-improving exchange. Therefore, applying Lemma 3, we get new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
    3. 2.3
      Agent $c_{1}$ $g$-champions $b_{1}$: Then we have $\langle(a_{1},c_{j}),(c_{1},b_{1}),(b_{1},a_{1})\rangle$ as a $\phi$-improving exchange. Thus, we get new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.∎
@@
We know that either $b_{1}$ or $c_{1}$ must envy $a_{1}$. Without loss of generality, assume $b_{1}$ envies $a_{1}$. Some leading agent must $g$-champion $b_{1}$. If $b_{1}$ self $g$-champions, then we can make Pareto improvement, thus we are done. If $a_{1}$ $g$-champions $b_{1}$, we have a $\phi$-improving exchange $\langle(a_{1},b_{1}),(b_{1},a_{1})\rangle$. Therefore, consider the case when agent $c_{1}$ $g$-champions $b_{1}$.
We know that some leading agent must $g$-champion $c_{1}$. If $c_{1}$ self $g$-champions, then we can make Pareto improvement. Suppose $b_{1}$ $g$-champion $c_{1}$. If either $c_{1}$ envies $b_{1}$ or $b_{1}$ envies $c_{1}$, then $(c_{1},b_{1},c_{1})$ is a Pareto improving cycle, thus we make a Pareto improvement. If not, we have $b_{1}$ and $c_{1}$ $g$-championing each other and $a_{1}$ envies neither $b_{1}$ nor $c_{1}$. We apply Lemma 5 to get Pareto improvement.
@@ 1. Case 1
    Agent $c_{1}$ picks a bundle $X_{b_{j}}$ from an agent $b_{j}\in\mathbb{B}$: Then, we construct a new allocation $Y$, by allocating $X_{b_{j}}$ to $c_{1}$, $T_{c_{1}}^{a_{1}}\cup g$ to agent $a_{1}$, and $X_{a_{1}}$ to $b_{j}$. Here, agent $a_{1}$ strictly improves. In this new allocation, no agent strongly envies any of $a_{1},b_{j}$ and $c_{1}$ as the bundles that these agents received are either $\mathsf{MES}$ bundles or complete bundles in the old allocation $X$. Agent $b_{j}$ does not strongly envy any agent as $v_{b}(Y_{b_{j}})\geq v_{b}(Y_{b_{1}})$ and agent $b_{1}$ does not strongly envy any agent. As $a_{1}$ received a strict improvement, she does not strongly envy any agent. Finally, we know that $c_{1}$ does not envy any agent from $\mathbb{B}\cup\mathbb{A}$ as it chose the best bundle. If $c_{1}$ strongly envies any agent in $\mathbb{C}$, we resolve them using Lemma 4. We now have $\phi(Y)>\phi(X)$.
@@ 3. Case 3
Therefore, Lemma 8 and Lemma 9 completes the proof of Theorem 1. Hence, EFX exists for three types of agents.
@@ 7 Conclusion
In this paper, we proved the existence of EFX allocations for instances with three types of agents, where agents of the same type have identical additive valuations. This result extends the prior work on EFX allocations for three agents [efx_3] and for two types of agents [mahara].
The techniques developed in this work may provide insights into more general settings. A natural direction for future research is to extend these results to agents with more general valuation functions, beyond the additive case. Another long standing open question is whether EFX allocations can be guaranteed for instances involving four or more agents.

### v3

@@ 49                  $U\leftarrow U\cup(X_{c_{i}}\setminus P)$
Algorithm 1 Iterated-Competition
@@
    Finally, if $P=X_{a_{i}}$ for some $a_{i}\in\mathbb{A}\setminus a_{1}$, then give $P$ to agent $b_{1}$ and $\{T_{b_{1}}^{a_{1}}\cup g\}$ to agent $a_{i}$. Resolve internal envies in $\mathbb{B}$ or $\mathbb{A}$, if any, using Lemma 3. This gives an EFX allocation $Y$ such that agent $a_{1}$ envies $b_{1}$. Therefore, by applying Lemma 8 on $Y$, we get an EFX allocation $Z$ such that $\phi(Z)>\phi(X)$.
@@
A virtual bundle appears in $O_{b}$ when Case 3Case 3a holds i.e., $\max_{b}O_{b}=X_{c_{i}}$ for some $i$ and $\max_{c}O_{c}=X_{b_{j}}$ for some $j\neq 1$, or Case 3(Case 3b)Case 3b(ii) holds, when $P=Q=X_{c_{i}}$ and either $b_{1}$ has a smaller winning subset than $c_{i}$ i.e. $|P_{b}|<|P_{c}|$ or both have $X_{c_{i}}$ as the winning subset i.e., $P_{b}=P_{c}=X_{c_{i}}$. When $P=\max_{b}O_{b}$ is a virtual bundle, then for any non-virtual i.e., real bundle $S\in O_{b}$, $P>_{b}S$ and hence $b_{1}$ does not envy the agent holding $S$. For any virtual bundle $S=X_{c_{i}}\setminus h\in O_{b}$, $S\neq P$, we have $P>_{b}X_{c_{i}}\setminus h$ for minimum valued item $h\in X_{c_{i}}$. Hence $b_{1}$ might envy the real bundle $X_{c_{i}}$ but it is not a strong envy. As $b_{j}$ and $b_{1}$ have the same valuation $v_{b}$, the lemma holds for any $b_{j}\in B$.∎
To prove the correctness of Case 3, we show that Invariant 1 is maintained as long as Algorithm 1 remains in Case 3, and that the algorithm exits from Case 3 in a finite number of iterations.
@@ 7 Agent $a_{1}$ is Not a Source in $E_{X}$
since $a_{1}$ is not a source, either $b_{1}$ or $c_{1}$ must envy $a_{1}$ by Proposition 4. Without loss of generality, assume $b_{1}$ envies $a_{1}$. We know that $a_{1}$ envies some agent from $\mathbb{B}\cup\mathbb{C}$. We proceed by considering the cases based on which agent is envied.
  1. Case 1
    Agent $a_{1}$ envies some agent $b_{j}\in\mathbb{B}$: Then, $\langle(a_{1},b_{j}),(b_{1},a_{1})\rangle$ is a pseudo-cycle. Therefore, using Lemma 4 we get a new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
  2. Case 2
    Agent $a_{1}$ envies some $c_{j}\in\mathbb{C}$: If $c_{1}$ envies $a_{1}$, then $\langle(a_{1},c_{j}),(c_{1},a_{1})\rangle$ is a pseudo-cycle. Thus we are done. Consider the case when $c_{1}$ does not envy $a_{1}$. We know that some leading agent must $g$-champion $b_{1}$. We consider all possible cases based on who $g$-champions $b_{1}$.
    1. Case 2.1
@@ 2. Case 2
    2. Case 2.2
      Agent $a_{1}$ $g$-champions $b_{1}$: Then $\langle(a_{1},b_{1}),(b_{1},a_{1})\rangle$ is a pseudo-cycle. Therefore, applying Lemma 4, we get new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
    3. Case 2.3
      Agent $c_{1}$ $g$-champions $b_{1}$: Then we have $\langle(a_{1},c_{j}),(c_{1},b_{1}),(b_{1},a_{1})\rangle$ as a pseudo-cycle. Thus, we get new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.∎
@@
We know that either $b_{1}$ or $c_{1}$ must envy $a_{1}$. Without loss of generality, assume $b_{1}$ envies $a_{1}$. Some leading agent must $g$-champion $b_{1}$. If $b_{1}$ self $g$-champions, then we can make a Pareto improvement, thus we are done. If $a_{1}$ $g$-champions $b_{1}$, we have a pseudo-cycle $\langle(a_{1},b_{1}),(b_{1},a_{1})\rangle$. Therefore, consider the case when agent $c_{1}$ $g$-champions $b_{1}$.
We know that some leading agent must $g$-champion $c_{1}$. If $c_{1}$ self $g$-champions, then we can make a Pareto improvement. Suppose $b_{1}$ $g$-champions $c_{1}$. If either $c_{1}$ envies $b_{1}$ or $b_{1}$ envies $c_{1}$, then $(c_{1},b_{1},c_{1})$ is a Pareto improving cycle, thus we make a Pareto improvement. If not, we have $b_{1}$ and $c_{1}$ $g$-championing each other and $a_{1}$ envies neither $b_{1}$ nor $c_{1}$. We apply Lemma 5 to get Pareto improvement.
@@ 1. Case 1
    Agent $c_{1}$ picks a bundle $X_{b_{j}}$ from an agent $b_{j}\in\mathbb{B}$: Then, we construct a new allocation $Y$, by allocating $X_{b_{j}}$ to $c_{1}$, $T_{c_{1}}^{a_{1}}\cup g$ to agent $a_{1}$, and $X_{a_{1}}$ to $b_{j}$. Here, agent $a_{1}$ strictly improves. In this new allocation, no agent strongly envies any of $a_{1},b_{j}$ and $c_{1}$ as the bundles that these agents received are either $\mathsf{MES}$ bundles or complete bundles in the old allocation $X$. Agent $b_{j}$ does not strongly envy any agent as $v_{b}(Y_{b_{j}})\geq v_{b}(Y_{b_{1}})$ and agent $b_{1}$ does not strongly envy any agent. As $a_{1}$ received a strict improvement, she does not strongly envy any agent. Finally, we know that $c_{1}$ does not envy any agent from $\mathbb{B}\cup\mathbb{A}$ as she chose the best bundle. If $c_{1}$ strongly envies any agent in $\mathbb{C}$, we resolve them using Lemma 3. We now have $\phi(Y)>\phi(X)$.
@@ 3. Case 3
Therefore, Corollary 5 and Corollary 6 complete the proof of Theorem 1. Hence, EFX exists for three types of agents.
@@ 8 Conclusion
In this paper, we proved the existence of EFX allocations for instances with three types of agents, where agents of the same type have identical additive valuations. Our result extends prior work on EFX allocations for three agents [efx_3], and for two types of agents [mahara].
The techniques developed in this work may provide insights into more general settings. A natural direction for future research is to extend these results to agents with more general valuation functions, beyond the additive case. Another long-standing open question is whether EFX allocations can be guaranteed for instances involving four or more agents.
@@
Acknowledgements
We are deeply grateful to the anonymous reviewers of ACM EC 2025 for their exceptionally thorough and thoughtful reviews. Their detailed feedback went far beyond the norm and significantly improved the clarity and quality of this work. Vishwa Prakash HV acknowledges support from the TCS-RSP Research Fellowship.